A population consists of the following five values: 13, 13, 12, 16, and 19.
  
(a)

List all samples of size 3, and compute the mean of each sample. (Round your Mean values to 2 decimal places.)

  
Sample Values Sum Mean
1 (Click to select)13,16,1913,13,1613,13,1213,12,16      
2 (Click to select)13,13,1213,13,1613,12,1613,16,19      
3 (Click to select)13,13,1613,13,1213,12,1613,13,19      
4 (Click to select)13,13,1213,16,1913,12,1613,13,16      
5 (Click to select)13,12,1613,12,1913,16,1913,13,12      
6 (Click to select)13,12,1613,16,1913,13,1213,12,19      
7 (Click to select)13,13,1213,16,1913,12,1612,16,19      
8 (Click to select)13,13,1213,12,1912,16,1913,16,19      
9 (Click to select)13,13,1213,12,1613,16,1912,16,19      
10 (Click to select)12,16,1913,13,1213,16,1913,12,19      

       

(b)

Compute the mean of the distribution of sample means and the population mean. (Round your answers to 2 decimal places.)

  
  
  Mean of the distribution of the sample mean   
  Population mean   
Answer #1
Sample values Sum mean
1 13,13,12 38 12.66667
2 13,13,16, 42 14
3 13,13,19 45 15
4 13,12,16 41 13.66667
5 13,12,19 44 14.66667
6 13,16,19 48 16
7 13,12,16 41 13.66667
8 13,12,19 44 14.66667
9 13,16,19 48 16
10 12,16,19 47 15.66667
Total 146

Mean of the distribution of the sample mean = 146 / 10 = 14.60

Population mean = (13 + 13 +12 + 16 + 19 ) / 5 = 14.60